OpenStudy (anonymous):

What is the solution set for |2x – 3| = 13?

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

Nope @Callie

OpenStudy (anonymous):

Hint : I2x-3I=13 IS 2x-3=13 and 2x-3=-13

OpenStudy (anonymous):

2 X 8-3 2X8+16 -3 +13 YES

OpenStudy (anonymous):

Solve that two equations

OpenStudy (anonymous):

@TAMATCHI123 got it?

OpenStudy (anonymous):

thankyou 8 and –8 5 and –5 8 and –5 –8 and 5

OpenStudy (anonymous):

What do u think?

OpenStudy (anonymous):

i would say it is 8 and -8

OpenStudy (anonymous):

is that the correct answer

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

2x-3=13 and 2x-3=-13

OpenStudy (anonymous):

Solve that two equations

OpenStudy (anonymous):

it would be 8 and -5

OpenStudy (anonymous):

YEP.....

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hey can you also help me on this one P = 2w + 2h for w. i really dont get it

OpenStudy (anonymous):

Take 2h to the left had side and divide both sides by 2

OpenStudy (anonymous):

hand side

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

do it first and show me what u got

OpenStudy (anonymous):

ok i got 1 for both sides

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i got 1

OpenStudy (anonymous):

p=2w+2h p-2h=2w w=(p-2h)/2

OpenStudy (anonymous):

wait so what is p here are my answers w = P – 2h w = P – 2h – 2 w = P - h / two w = P - 2h / two

OpenStudy (anonymous):

w = P - 2h / two

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

yes that is right